# DualMAR: Medical-Augmented Representation from Dual-Expertise Perspectives

### v1

@@ 4.2 Prediction Results
Table 2 demonstrates that DualMAR consistently outperforms existing baselines across both diagnosis and HF prediction tasks for MIMIC-III dataset. The proposed framework consistently performs the best, underlining the efficacy of external priors from clinical KGs for prediction.
Table 2: Diagnosis and Heart Failure Prediction.
Diagnosis Prediction HF Prediction
Models w-$\mathbf{F_{1}}$ R@10 R@20 AUC $\mathbf{F_{1}}$
Dipole 14.66(0.2) 28.73(0.2) 29.44(0.2) 82.08(0.3) 70.35(0.2)
Deepr 11.68(0.1) 26.47(0.1) 27.53(0.1) 81.36(0.1) 69.54(0.1)
RETAIN 18.37(0.8) 32.12(0.8) 32.54(0.6) 83.21(0.3) 71.32(0.2)
Timeline 20.46(0.2) 30.73(0.1) 34.83(0.1) 82.34(0.3) 71.03(0.2)
GRAM 20.78(0.1) 34.17(0.2) 35.46(0.3) 81.55(0.2) 68.78(0.1)
KAME 21.10(0.1) 29.97(0.2) 33.99(0.2) 82.88(0.1) 72.03(0.1)
CGL 22.63(0.2) 33.64(0.3) 37.87(0.2) 84.19(0.2) 71.77(0.1)
G-BERT 22.28(0.3) 35.62(0.2) 36.46(0.2) 81.50(0.2) 71.18(0.1)
HiTANet 23.15(0.2) 34.68(0.3) 35.97(0.1) 85.13(0.1) 73.15(0.2)
THAM 23.92(0.1) 38.64(0.2) 39.97(0.1) 83.24(0.1) 74.12(0.1)
DualMAR 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.2})}$ 41.86${}_{(\textbf{0.3})}$ 86.53${}_{(\textbf{0.1})}$ 75.35${}_{(\textbf{0.1})}$
Table 3: Real-time and Non-chronic Diagnosis Prediction.
@@
THAM 24.52(0.4) 40.31(0.2) 41.45(0.3) 19.93(0.3) 33.53(0.4)
DualMAR 26.89${}_{(\textbf{0.3})}$ 42.86${}_{(\textbf{0.1})}$ 43.15${}_{(\textbf{0.2})}$ 21.69${}_{(\textbf{0.2})}$ 34.35${}_{(\textbf{0.4})}$
Prediction Results on MIMIC-III. We report the average performance (%) and standard deviation (in bracket) of each model over 10 runs. The best results for each metrics are highlighted.
4.2.1 Real-time & Non-chronic Diagnosis Prediction.
Table 3 shows the results of two supplementary experiments, real-time and non-chronic diagnosis predictions, respectively.
@@
    Real-time (RT) Diagnosis: To fairly compare performance of all baselines, we feed lab tests information for all baselines and proposed model, which not only augments representation for diagnosis prediction but also mimics real-time diagnosis of physician within the beginning of an admission. We can observe that the proposed model still outperforms all baselines, which has even a larger improvement on w-F1 over regular diagnosis. We suppose that DualMAR, which originally relies on lab tests for auxiliary diagnosis, will be more suitable for lab features input than other “uni-aspect” baselines.
@@
    Non-chronic (NC) Diagnosis: Since MIMIC-III is mostly ICU data, we remove all chronic diseases in predictions and focus on predicting acute condition codes for future visits. We can observe that DualMAR still achieves superior performance, but all baselines and proposed model get relatively low values of metrics compared to regular diagnosis. Note that, R@20 will be omitted in report since the average number of diagnoses is less than 10 within each admission. We conjecture that the removal of chronic diseases in admissions skews the distribution, making rare diseases more prominent and, consequently, making the prediction task more challenging.
@@ 4.3 Ablation Study
To evaluate the effectiveness of involved modules in DualMAR, we compare performance of variants with some modules removed or replaced. We use both future diagnosis and HF prediction tasks on MIMIC-III as examples:
  1. 1.
    DualMARa: Removing Diagnosis KG and KG embedding modules, initializing embeddings randomly.
  2. 2.
    DualMARb: Replacing the embedding of the Diagnosis KG with GloVe, as used in GRAM Choi et al. (2017), to only focus on hierarchical information.
  3. 3.
    DualMARc: Replacing the embedding of the Diagnosis KG with text-embedding-3-large OpenAI (2024) to only retrieve semantic information.
  4. 4.
    DualMARd: Removing the Graph Learning module and directly feeding KG embeddings into the Attention layer.
  5. 5.
    DualMARe: Replacing EHR graph by a simplified graph with disease complications, excluding lab test nodes.
  6. 6.
    DualMARf: Replacing Attention by GRU modules.
  7. 7.
    DualMARg: Removing the proxy-task learning module and training the model directly on the downstream task.
Table 4: Ablation Results of the Variants of DualMAR.
@@
DualMARa 21.21(0.2) 35.25(0.3) 36.46(0.3) 80.75(0.4) 69.81(0.3)
DualMARb 22.32(0.3) 36.28(0.3) 36.95(0.4) 81.66(0.3) 70.01(0.3)
DualMARc 23.15(0.3) 37.78(0.4) 38.79(0.4) 83.16(0.3) 72.73(0.3)
DualMARd 22.49(0.3) 34.86(0.3) 35.79(0.4) 80.90(0.3) 69.24(0.3)
DualMARe 24.58(0.3) 39.78(0.4) 40.46(0.4) 83.92(0.3) 73.62(0.3)
DualMARf 24.96(0.3) 40.16(0.3) 40.87(0.4) 85.81(0.3) 73.96(0.3)
DualMARg 23.74(0.3) 38.19(0.3) 39.52(0.4) 84.73(0.3) 72.85(0.3)
DualMAR 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.3})}$ 41.86${}_{(\textbf{0.3})}$ 86.53${}_{(\textbf{0.3})}$ 75.35${}_{(\textbf{0.3})}$
- no lab 24.21(0.3) 39.47(0.3) 40.35(0.3) 85.73(0.3) 73.92(0.3)
We report the results of DualMAR and its variants DualMARa-g in Table 4. The full DualMAR model consistently outperforms all variants, proving the effectiveness of our framework The comparisons between DualMAR and DualMARb-c highlight the importance of both hierarchical and semantic information from the KG, even though each alone can enhance predictions based on DualMARa. Additionally, the graph learning DualMARd and attention modules DualMARf are critical for refining representations, with the inclusion of lab tests further enhancing model performance in terms of DualMARe. The significant improvement of DualMAR over DualMARg underscores the essential role of the proxy-task learning module.
Notably, for cases without lab test data, the No LAB model, which uses modified GNN and skips pretraining, still achieves top-tier performance among baselines, demonstrating the effectiveness and generalizability of DualMAR.
@@ 4.4 Case Study
4.4.1 Less-frequent Diagnosis Prediction
We also design experiments focusing on predicting less frequent condition codes in MIMIC-III, similar to rare ICD code prediction tasks. We select and predict 100 ICD codes that occur fewer than 10 times in the dataset. We follow the previous settings which excludes records without rare codes to stabilize training. We use recall at 5/8/15/20 for evaluation. Figure 2(a) compares DualMAR with baselines using ontology or transformer architectures. DualMAR outperforms other selected methods on R@5, R@8, R@15, and R@20. Notably, CGL and G-BERT show significant improvements, indicating that external knowledge bases can aid in diagnosing rare conditions. This underscores the importance of incorporating medical ontologies or complex KGs to enhance model resilience against EHR data limitations.
@@ 4.4.1 Less-frequent Diagnosis Prediction
Figure 2: Results for the $1^{st}$ and $2^{nd}$ case studies, full results are shown in Appendix F.
4.4.2 Advanced Analysis of KG Embeddings
To assess the importance of each component in our Diagnosis KG, we divided it into three parts: ICD9-Hierarchy KGa, Ontology-KG KGb, and LLM-KG KGc. Note that, the reason why both KGa+b and KGa+b+c are ignored in experiment is that knowledge in KGa is mostly covered by KGb, so that they will get same results as other existing settings. We then evaluated the predictive performance of the model with these varying complexities. Figure 2(b) shows how the complexity and embedding dimensions of KGs influence diagnosis prediction. Key observations include: (1) LLM-generated triples (KGc) enhance comprehensive KG representation, as seen in KGa+c and KGb+c. (2) Ontology-KG (KGb) from diverse data sources is more reliable for accurate predictions than ICD9-Hierarchy (KGa). (3) Embedding dimensions significantly impact predictions; higher dimensions offer a deeper understanding but can lead to overfitting, especially with simpler KGs like KGa.
@@ 4.4.3 Accessibility
We evaluated the accessibility of DualMAR with large-scale KGs, comparing it to GraphCare Jiang et al. (2023) using an NVIDIA RTX-4090 GPU with 24GB of memory. For a fair comparison, we adapted UMLS-GPT-KG in GraphCare to our Diagnosis KG, adjusting only the batch size due to its end-to-end structure without pretraining module. Table 5 shows average GPU usage for diagnosis prediction. GraphCare struggles with training on common commercial GPUs when transitioning from CCS KG to ICD-9 KG, even with small batch sizes, limiting its scalability with large KGs. In contrast, DualMAR efficiently separates learning processes for KGs and EHR data, allowing it to scale to larger KGs and making it more deployable across various EHR tasks and datasets.
@@
DualMAR - batch sizes: 256, 32, and 32
@@ 5 Conclusion
In this study, we introduced DualMAR, a framework designed to enhance EHR prediction tasks by leveraging both individual health records and public knowledge bases. Through the construction of a bi-hierarchical diagnosis KG augmented by LLMs and the design of a proxy-task learning via lab results, DualMAR significantly improves the representation of patient embeddings and disease-related predictions. Our experiments demonstrate that DualMAR outperforms existing state-of-the-art models, highlighting the potential of integrating comprehensive hierarchical and semantic information from KGs into EHR models. This approach not only broadens the scope of medical-domain knowledge but also addresses the limitations of traditional EHR models by incorporating more complex and nuanced medical data. Future work could explore expanding DualMAR to adapt to large-scale KGs and apply it to a broader range of healthcare prediction tasks beyond diagnostics. Additionally, this work highlights the crucial role of laboratory data in clinical decision-making, suggesting that further investigation into leveraging lab tests as auxiliary inputs for both diagnosis and treatment predictions remains a promising direction.
@@
Appendix A KG Construction
A.1 Data Sources in Healthcare KG.
To develop a comprehensive KG for studying diseases, we consider two primary resources related to diseases and phenotypes, and incorporate three additional resources to enhance the connectivity of our Diagnosis KG. Notably, these data resources—whether widely-used standardized ontologies or direct readouts from experimental measurements—primarily focus on extensive coverage across disease, phenotype, and drug entities. We begin with a brief introduction of primary data resources used to construct component of the Diagnosis KG:
  1. 1.
    DrugBank. DrugBank Knox et al. [2024] is a comprehensive resource containing detailed pharmaceutical knowledge. We retrieved the latest version (5.1.12), published on March 14, 2024. Our focus is on synergistic drug interactions, which represent the bidirectional connections between two drugs.
  2. 2.
    DrugCentral. DrugCentral Ursu et al. [2016] is a curated database that provides information on drug-disease interactions, including indications, contraindications, and off-label uses. We utilized the updated SQL Database, released on November 1, 2023, for our study.
  3. 3.
    HPO. The Human Phenotype Ontology Gargano et al. [2024] offers detailed information on phenotype abnormalities associated with diseases. We used one of the most recent updates, from April 19, 2024, focusing on disease-phenotype and phenotype-phenotype relationships, and extracted verified associations with expertly curated annotations.
  4. 4.
    ICD-9-CM Disease Ontology. ICD-9-CM Organization et al. [1988] is a coding system used by health insurers to classify medical conditions for billing purposes, and it underlies the diagnostic information in most EHR data. For our purposes, we extracted the parent-child relationships among codes to describe disease interactions.
  5. 5.
    SIDER. The Side Effect Resource Kuhn et al. [2016] contains data on side-effect phenotype caused by various drugs. We retrieved this information from the data release dated October 21, 2015.
A.2 Data Sources in LLM-generated KG.
We also extract textual features for disease nodes in the KG from several widely-used clinical knowledge bases. The in-context learning capabilities of LLMs are employed to convert this textual information into triples, which are then integrated into the KG. The following four public data sources are used for triple generation:
  1. 1.
    Mayo Clinic. Mayo Clinic Brennan et al. [1998] is a nonprofit academic medical center and biomedical research institution that maintains a knowledge base with clinical information on over 2000 diseases. We collected Mayo Clinic web knowledge in 2023 and mainly focus on both “Symptoms & Causes” (Overview, Symptoms, When to see a doctor, Causes, Risk Factors, Complications, and Prevention) and “Diagnosis & Treatment” parts for each condition.
  2. 2.
    Orphanet. Orphanet Weinreich et al. [2008] is a database dedicated to rare diseases. We collected data in 2023, focusing on information about definitions, prevalence, treatment, and clinical descriptions for rare diseases.
  3. 3.
    Rare Disease Database. Rare Disease Database Schieppati et al. [2008], created by a nonprofit organization NORD, serves as a supplementary source for rare diseases. In 2023, we collected data to retrieve information on disease overviews, symptoms, causes, treatments, and clinical trials.
  4. 4.
    Wikipedia. Wikipedia is a free, open-content online encyclopedia maintained by volunteers worldwide. It includes an official list of ICD-9 condition codes. We web-scraped the overall hierarchy of ICD-9 codes and gathered related information for diseases with individual Wikipedia pages through valid hyperlinks. For example, the page for “Typhoid Fever” provides detailed information on symptoms, causes, diagnosis, prevention, treatment, epidemiology, history, terminology, and societal impact, offering more comprehensive and diverse knowledge compared to other textual sources.
Name Prompt Template
A.3 Prompting LLM
It is important to note that since all textual information for disease nodes is sourced from either publicly verified or expertly curated data sources, issues of data quality or hallucination in the medical knowledge generated by the LLM are not a primary concern. In this section, we focus on how to utilize the in-context learning capabilities of LLMs to transform this textual information into structured triples that can be directly integrated into the KG.
The LLM used in this study is Llama-3-8b. Table 6 illustrates a carefully designed prompt for constructing triples from disease conditions. The primary objective is to harness the extensive knowledge embedded in the LLM to extract meaningful triples consisting of two entities and a relationship. Llama-3-8b was chosen for inference in this study for several reasons: first, its relatively small parameter size makes it easier to deploy and use; second, it is one of the latest and most widely adopted open-source models; and third, the task at hand is common and straightforward, not requiring the large parameter size or fine-tuning that other medical-domain LLMs might offer.
Our strategy begins with a prompt related to a specific medical condition. The LLM is then tasked with generating a list of triples, following the format [ENTITY 1, RELATIONSHIP, ENTITY 2], where both ENTITY 1 and ENTITY 2 are nouns. The goal is to generate these triples with both breadth (capturing potential abbreviations or partial matches) and accuracy (ensuring that each triple is based on existing information).
Although triple generation is relatively straightforward for LLMs, it still remains two issues:
  • Sometimes, the output is either empty or consists of garbled text, which we consider as deprecated results (approximately 8% of cases).
  • Occasionally, the output triples are incomplete or end with “…” (approximately 3% of cases).
To mitigate these issues, we employ iterated running and re-read prompting strategies to ensure stable outputs. For iterated running, we set the hyperparameter $x$ to control the number of times the entire textual data is re-prompted, with $x=2$ in this study. For re-read prompting, we set the hyperparameter $y$ to repeatedly feed the same prompt into the model, allowing the LLM to reconsider and refine its answers, with $y=1$ in this study. These strategies help the LLM generate comprehensive triples from all textual information, with the final output curated through manual checks.
Considering accessibility, we chose to utilize open-source LLMs for the triple generation step, as they are more accessible and convenient to deploy on local servers compared to the GPT series. In this experiment, while Llama-3-8b serves as our inference LLM example, other widely-used LLMs like Llama-2-13b or fine-tuned models like PMC-Llama can also be used in this process.
Appendix B KG Embeddings Comparison
Most current hierarchical KG embeddings are designed to capture and represent hierarchy information in KGs using various geometric structures, including Euclidean and non-Euclidean geometries. Euclidean-based methods typically focus on projections across different coordinates to distinguish entities at various levels. Compared to non-Euclidean approaches like hyperbolic and spherical embeddings, Euclidean projections can efficiently retrieve hierarchical information without additional processes or high computational costs. Additionally, we can separately train embeddings through link prediction tasks, allowing GPU usage to be decoupled from the graph learning process, ensuring stable performance even with more complex KGs. We propose projecting each entity in the KG onto polar coordinates, which enables more efficient embedding retrieval.
Table 7 compares the polar-space-based KG embedding method, which is an adjusted embedding method refered from HAKE Zhang et al. [2020], with five widely-used baselines Bordes et al. [2013], Sun et al. [2019], Nickel and Kiela [2017], Fionda and Pirrò [2020], demonstrating the advantages of our approach in retrieving both hierarchical and semantic features from the KG. Performance is evaluated using MRR and HITS@k metrics for regular link prediction tasks, with k set to 1, 3, and 10 in our experiments.

### v2

@@ 4.2 Prediction Results
Table 3: Prediction Results on MIMIC-III and MIMIC-IV for Diagnosis (Task 1) and HF (Task 2) Prediction. We report the average performance (%) and standard deviation (in bracket) of each model over 10 runs. The best results are highlighted.
Model Task 1: Diagnosis Prediction Task 2: Heart Failure Prediction
MIMIC-III MIMIC-IV MIMIC-III MIMIC-IV
w-F1 R@10 R@20 w-F1 R@10 R@20 AUC F1 AUC F1
RETAIN 18.37(0.8) 32.12(0.8) 32.54(0.6) 23.11(0.8) 37.32(0.8) 40.15(0.6) 83.21(0.3) 71.32(0.2) 84.14(0.3) 71.23(0.2)
Timeline 20.46(0.2) 30.73(0.1) 34.83(0.1) 23.76(0.2) 37.89(0.1) 40.87(0.1) 82.34(0.3) 71.03(0.2) 83.45(0.3) 72.30(0.2)
GCT 20.66(0.2) 32.73(0.2) 35.44(0.2) 24.16(0.2) 38.24(0.2) 41.65(0.2) 82.08(0.3) 70.35(0.2) 84.80(0.3) 69.52(0.2)
CGL 22.63(0.2) 33.64(0.3) 37.87(0.2) 25.74(0.2) 39.23(0.3) 42.67(0.2) 84.19(0.2) 71.77(0.1) 87.91(0.2) 70.71(0.3)
SeqCare 21.78(0.1) 34.17(0.2) 36.46(0.3) 24.39(0.1) 38.42(0.2) 41.62(0.3) 81.55(0.2) 68.78(0.1) 85.55(0.2) 69.82(0.4)
RAM-EHR 23.71(0.1) 37.97(0.2) 40.18(0.2) 27.01(0.1) 42.86(0.2) 46.92(0.2) 82.88(0.1) 72.03(0.1) 84.80(0.2) 72.34(0.5)
G-BERT 22.28(0.3) 35.62(0.2) 36.46(0.2) 25.12(0.3) 39.91(0.2) 43.25(0.2) 81.50(0.2) 71.18(0.1) 85.76(0.2) 72.88(0.1)
HiTANet 23.15(0.2) 34.68(0.3) 35.97(0.1) 24.53(0.2) 38.42(0.3) 41.89(0.1) 85.13(0.1) 73.15(0.2) 86.34(0.4) 71.35(0.2)
Med-BERT 21.68(0.1) 33.47(0.1) 36.53(0.1) 23.58(0.1) 36.79(0.1) 40.45(0.1) 81.36(0.1) 69.54(0.1) 83.61(0.5) 70.46(0.1)
GT-BEHRT 23.92(0.1) 38.64(0.2) 39.97(0.1) 26.97(0.3) 43.07(0.4) 47.19(0.2) 83.24(0.1) 74.12(0.1) 87.43(0.3) 72.26(0.2)
DuaLK 25.37(0.3) 40.52(0.2) 41.86(0.3) 29.87(0.3) 45.66(0.2) 51.73(0.3) 86.53(0.1) 75.35(0.1) 90.32(0.2) 73.54(0.1)
Table 4: Supplementary Prediction Results on MIMIC-III for Real-time and Non-chronic Diagnosis Prediction.
@@
Med-BERT 22.74(0.3) 35.27(0.4) 36.54(0.4) 17.94(0.3) 30.84(0.4)
GT-BEHRT 24.53(0.4) 40.31(0.2) 41.45(0.3) 19.93(0.3) 33.53(0.4)
DuaLK 26.89${}_{(\textbf{0.3})}$ 42.86${}_{(\textbf{0.1})}$ 43.15${}_{(\textbf{0.2})}$ 21.69${}_{(\textbf{0.2})}$ 34.35${}_{(\textbf{0.4})}$
Table 3 demonstrates that DuaLK consistently outperforms existing baselines across both diagnosis and heart failure prediction tasks on both MIMIC-III and MIMIC-IV datasets, which underlines the efficacy of external priors from clinical KGs for prediction. For diagnosis prediction, we can observe that DuaLK boosts the weighted F1 score by approximately 6% relative to the best baseline on MIMIC-III. Similar trends can be observed in the R@10 and R@20 metrics, where DuaLK surpasses the strongest baselines by around 4–5%. The improvements on MIMIC-IV are even more pronounced, since DuaLK reaches a w-$F_{1}$ of 29.87%, an improvement of nearly 11%. The R@10 and R@20 metrics follow suit, with DuaLK outperforming the best baseline by about 6% and 10%, respectively. Regarding the HF prediction task, the absolute gains are somewhat smaller but remain consistent.
Table 4 shows the results of two modified tasks, real-time and non-chronic diagnoses, on the MIMIC-III dataset:
@@
    Real-time (RT) Diagnosis: To fairly compare performance of all baselines, we feed lab tests information for all baselines and proposed model, which not only augments representation for diagnosis prediction but also mimics real-time diagnosis of physician within the beginning of an admission. We can observe that the proposed model still outperforms all baselines, which has even a larger improvement on w-F1 over regular diagnosis. We suppose that DuaLK, which originally relies on lab tests for auxiliary diagnosis, will be more suitable for lab features input than other baselines.
@@
    Non-chronic (NC) Diagnosis: Since MIMIC-III mostly collects ICU data, we remove all chronic diseases and focus on predicting acute diseases for future visits. Specifically, there are 142 codes removed by the chronic condition indicator  [^2] https://hcup-us.ahrq.gov/toolssoftware/chronic/chronic.jsp for ICD-9-CM, and most of them frequently appears in admission records. Therefore, all models suffer performance degradations compared to the regular setting. We can observe that DuaLK still achieves superior performance, but all baselines and proposed model get relatively low values of metrics compared to regular diagnosis. We conjecture that the removal of chronic diseases in admissions skews the distribution, making rare diseases more prominent and, consequently, making the task more challenging.
Overall, these results suggest that the dual integration of public knowledge from clinical knowledge graphs and individualized patient data is crucial. By leveraging both external priors and patient-specific nuances, DuaLK not only bridges the gap between structured medical information and dynamic EHR data but also consistently outperforms models that rely on either temporal learning, medical ontology, or Transformer-based architectures alone.
@@ 4.3 Ablation Study
Table 5: Ablation Results of the Variants of DuaLK.
@@
DuaLKa 21.21(0.2) 35.25(0.3) 36.46(0.3) 80.75(0.4) 69.81(0.3)
DuaLKb 22.32(0.3) 36.28(0.3) 36.95(0.4) 81.66(0.3) 70.01(0.3)
DuaLKc 23.15(0.3) 37.78(0.4) 38.79(0.4) 83.16(0.3) 72.73(0.3)
DuaLKd 22.49(0.3) 34.86(0.3) 35.79(0.4) 80.90(0.3) 69.24(0.3)
DuaLKe 24.58(0.3) 39.78(0.4) 40.46(0.4) 83.92(0.3) 73.62(0.3)
DuaLKf 24.96(0.3) 40.16(0.3) 40.87(0.4) 85.81(0.3) 73.96(0.3)
PrimeKGg 23.74(0.3) 38.19(0.3) 39.52(0.4) 84.73(0.3) 72.85(0.3)
GraphCareh 24.14(-) 39.89(-) 40.14(-) 83.73(-) 76.28(-)
DuaLK 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.3})}$ 41.86${}_{(\textbf{0.3})}$ 86.53${}_{(\textbf{0.3})}$ 75.35${}_{(\textbf{0.3})}$
- w/o lab 24.21(0.3) 39.47(0.3) 40.35(0.3) 85.73(0.3) 73.92(0.3)
To evaluate the effectiveness of involved modules in DuaLK, we compare performance of variants with some modules removed or replaced. We use both future diagnosis and HF prediction tasks on MIMIC-III as examples:
  1. 1.
    DuaLKa: Replacing the embedding of the Diagnosis KG with GloVe, as used in GRAM [7], to only focus on hierarchical information for diseases.
  2. 2.
    DuaLKb: Replacing the embedding of the Diagnosis KG with text-embedding-3-large [52] to only retrieve semantic information for diseases.
  3. 3.
    DuaLKc: Replacing the link prediction tasks with minimizing distances among connected codes and maximizing distances among unconnected ones.
  4. 4.
    DuaLKd: Removing the Graph Learning module and directly feeding KG embeddings into the Attention layer.
  5. 5.
    DuaLKe: Replacing EHR graph by a simplified graph with disease complications, excluding lab test nodes.
  6. 6.
    DuaLKf: Replacing Attention by GRU modules.
  7. 7.
    PrimeKGg: Using PrimeKG as the KG for initialization of disease embeddings by mapping identifier from MONDO to ICD-9-CM.
  8. 8.
    GraphCareh: Feeding learned embeddings from the Diagnosis KG into the predictor used by Graphcare aligning to the same encoding system.
The results of ablation variants are shown in Table 5. DuaLK consistently outperforms its variants, validating the effectiveness of each module. The comparison with DuaLKa-b shows that both hierarchical and semantic information in the KG contribute to performance, even though each alone provides some benefit. Results from DuaLKd-f demonstrate the importance of graph learning and attention for representation refinement, with lab tests further enhancing performance. The substantial gain over PrimeKGg highlights the value of a task-specific KG with relevant entities. Although GraphCareh achieves higher F1, its higher computational cost underscores the efficiency of leveraging multi-aspect information beyond fine-grained graphs. The model w/o lab using modified GNN and pretraining tasks in Sherbet [5], still achieves superiority over baselines, demonstrating the generalizability of DuaLK.
Figure 3: Performance by EHR Training Data Sizes. Values on the x-axis indicate % of the entire training data. The dotted lines separate two ranges: [1, 10] and [10, 100] (%).
@@ 4.4 Case Study
4.4.1 Effect of EHR Data Size
By incorporating the diagnosis KG as medical prior knowledge before the training phase, DuaLK is expected to exhibit greater robustness to variations in training data size. Hence, it is crucial to evaluate how well models can provide accurate diagnoses with limited labeled data, both with and without prior knowledge. Figure 3 presents the results for R@10 in Diagnosis Prediction and AUC in heart failure prediction. As shown, all models experience performance degradation as the amount of labeled data decreases, especially when less than 10% of the training set is available. However, DuaLK consistently outperforms other baselines, demonstrating that knowledge injection from the KG effectively enhances information utilization, even in scenarios with scarce labeled data.
4.4.2 Diagnosis for Low-Prevalence Diseases
We also design experiments focusing on predicting less frequent condition codes in MIMIC-III, similar to low-prevalence medical coding tasks. We select and predict 100 ICD codes that occur fewer than 10 times in the dataset. We follow the previous settings which excludes records without low-prevalence diseases to stabilize training. We use recall at 5/8/15/20 for evaluation. Figure 4(a) compares DuaLK with baselines using ontology or transformer architectures. DuaLK outperforms other selected methods on R@5, R@8, R@15, and R@20. Notably, CGL and G-BERT show significant improvements, indicating that external knowledge bases can aid in diagnosing less-frequent conditions. This underscores the importance of incorporating medical ontologies or complex KGs to enhance model resilience against EHR data limitations.
4.4.3 Advanced Analysis of KG Embeddings
To assess the importance of each component in our Diagnosis KG, we divided it into three parts: ICD9-Hierarchy KGa, Ontology-KG KGb, and LLM-KG KGc. We then evaluated the predictive performance of the model with these varying complexities. Figure 4(b) shows how the complexity and embedding dimensions of KGs influence diagnosis prediction. Key observations include: (1) LLM-generated triples (KGc) enhance comprehensive KG representation, as seen in KGa+c and KGb+c. (2) Ontology-KG (KGb) from diverse data sources is more reliable for accurate predictions than ICD9-Hierarchy (KGa). (3) Embedding dimensions significantly impact predictions; higher dimensions offer a deeper understanding but can lead to overfitting, especially with simpler KGs like KGa.
@@ 4.4.3 Advanced Analysis of KG Embeddings
Figure 4: Results for the $1^{st}$ and $2^{nd}$ case studies, full results are shown in Appendix 5.3.
@@ 4.4.4 Accessibility
We evaluated the accessibility of DuaLK with large-scale KGs, comparing it to GraphCare [9] using an NVIDIA RTX-4090 GPU with 24GB of memory. For a fair comparison, we adapted UMLS-GPT-KG in GraphCare to our Diagnosis KG, adjusting only the batch size due to its end-to-end structure without pretraining module. Table 6 shows average GPU usage for diagnosis prediction. GraphCare struggles with training on common commercial GPUs when transitioning from CCS KG to ICD-9 KG, even with small batch sizes, limiting its scalability with large KGs. In contrast, DuaLK efficiently separates learning processes for KGs and EHR data, allowing it to scale to larger KGs and making it more deployable across various EHR tasks and datasets.
@@
DuaLK - batch sizes: 256, 32, and 32
@@ 5 Conclusion
We proposed DuaLK, a dual-expertise framework for enhancing clinical prediction by jointly leveraging individual lab results and structured medical knowledge. DuaLK generates a bi-hierarchical Diagnosis KG enriched via LLM-generated triples, capturing both semantic and hierarchical information essential for clinical prediction. Moreover, a lab-informed proxy-task learning aligns the model with real-world diagnostic processes by guiding stepwise clinical reasoning for diagnosis. These two perspectives are then integrated via a encoder-decoder architecture, enabling more precise and interpretable predictions. Experimental results demonstrate that DuaLK consistently outperforms existing predictive baselines across multiple clinical tasks, which highlights the potential of combining structured knowledge and clinical evidence to advance EHR modeling. Overall, DuaLK not only broadens the scope of medical-domain knowledge but also addresses the limitations of traditional EHR models by incorporating more complex and nuanced medical data. Future work could explore expanding DuaLK to adapt to large-scale KGs and apply it to a broader range of healthcare prediction tasks beyond diagnostics. Additionally, this work highlights the crucial role of laboratory data in clinical decision-making, suggesting that further investigation into leveraging lab tests as auxiliary inputs for both diagnosis and treatment predictions remains a promising direction.
[…]
@@
5.1 KG Embeddings Comparison